# Bongard in Wonderland: Visual Puzzles that Still Make AI Go Mad?

### v2

@@ 1 Introduction
Visual reasoning, the ability to understand, interpret, and reason about the visual world, is a fundamental aspect of human intelligence (Marr, 2010). It allows us to navigate our environment, interact with objects, and make sense of complex visual scenes. In recent years, the field of artificial intelligence (AI) has advanced rapidly toward replicating aspects of this visual reasoning, with significant focus placed on Vision-Language Models (VLMs) (Bordes et al., 2024; Liu et al., 2023, 2024). These models integrate visual and textual information to generate descriptive content, aiming to mimic how humans comprehend and reason about the physical world. Because of their human-like responses, VLMs often create the illusion of possessing human-like perception and intelligence. However, as recent work shows, VLMs and the Large Language Models (LLM) on which they are based have dramatic shortcomings in the case of reasoning (Nezhurina et al., 2024) and visual perception (Kamath et al., 2023; Rahmanzadehgervi et al., 2024; Geigle et al., 2024; Gou et al., 2024) or their combination (Zhou et al., 2023; Zhang et al., 2024; Wang et al., 2024b).
@@
While traditional machine learning approaches have made some early progress on BPs (Raghuraman et al., 2023; Depeweg et al., 2024), the potential of VLMs remains largely unexplored. Given that VLMs already struggle with recognizing relatively simple visual patterns (Rahmanzadehgervi et al., 2024; Zhang et al., 2024), BPs present a particularly challenging task, offering a valuable framework for investigating which patterns are easier or harder for state-of-the-art models to identify. Recent work has explored BPs in the context of VLM evaluation, but in a more limited capacity (Małkiński et al., 2024). While their study provides meaningful insights, it only considers a classification-based evaluation and does not analyze model behavior in greater depth. This highlights the need for a more comprehensive investigation into how VLMs process and reason about BPs.
@@ 2 Related Work
Benchmarks for VLMs. Benchmarks specifically designed for VLMs usually involve complex tasks such as image captioning, scene or diagram understanding, visual question answering (VQA), or visual-commonsense reasoning (Masry et al., 2022; Yue et al., 2024; Rahmanzadehgervi et al., 2024; Kamath et al., 2023; Antol et al., 2015; Johnson et al., 2017; Hong et al., 2021; Zellers et al., 2019; Hudson & Manning, 2019; Yuxuan et al., 2024; Yuan et al., 2023). Yet, most of these only require simple reasoning abilities. More recent benchmarks have been introduced to probe advanced reasoning skills, e.g., logical learning (Helff et al., 2023; Vedantam et al., 2021; Xiao et al., 2024), mathematical reasoning (Lu et al., 2024) or analogy-based visual reasoning (Chollet, 2019; Moskvichev et al., 2023; Zhang et al., 2019). Małkiński et al., 2024 considered BPs for evaluating VLMs and developed a benchmark inspired by them. While they provide meaningful insights, they only consider a classification setting for the evaluation and do not investigate the model behaviour in more depth. Generally, the shift towards more cognitively demanding tasks is promising, comprehensive diagnostic evaluations of VLMs’ reasoning capabilities that pinpoint sources of error and model limitations remain scarce. Furthermore, the degree to which these models genuinely comprehend complex, abstract visual concepts is yet to be fully investigated.
Open-Ended Visual Reasoning. Most existing benchmarks rely on classification tasks (Chollet, 2019; Helff et al., 2023; Vedantam et al., 2021), multiple-choice formats (Zhang et al., 2019; Lu et al., 2024), or synthetic environments with a limited set of operations (Moskvichev et al., 2023). While these setups tackle pattern recognition and visual reasoning, they do not fully capture real-world complexity. Open-ended visual puzzles, such as BPs, present a greater challenge by requiring models to map visual input to text without predefined rules or language biases. This lack of structured guidance makes the task more demanding, better reflecting real-world reasoning.
@@ 3 Bongard Problems and VLMs
To achieve this, we designed a prompt for each BP to specifically address the ground truth concepts. For any given image among the 12 images of a BP, the prompt asks whether the concept from the left or the right side of the BP applies to that image. Given two example prompts for BP#16 and BP#55 (Suppl.  A.1 and Suppl.  A.1) and ground truth rules [^2] https://www.foundalis.com/res/bps/bongard_problems_solutions.htm, the prompts for each BP get auto-generated by an LLM (cf. Suppl.  A.1).
@@ 4 Experimental Evaluation
Our experimental evaluations investigate to what extent state-of-the-art VLMs can solve Bongard problems. For this, we first evaluate the models quantitatively on all 100 puzzles based on the settings of Task 1 and Task 1.1. We then compare their performance against humans and investigate them qualitatively in more detail based on the settings of Task 2 and Task 3. We hereby address the following research questions: (Q1: Task 1 + Task 1.1) How well can state-of-the-art VLMs solve Bongard problems? (Q2) How do VLMs perform in comparison to humans? (Q3: Task 2) How accurately can VLMs classify images of the BPs given the ground truth rule? (Q4: Task 3) How accurately can VLMs generate an underlying hypothesis?
Models. We evaluated the proprietary models o1 (OpenAI, 2024b), GPT-4o (OpenAI, 2024a), Claude 3.5 Sonnet (Antropic, 2024), Gemini 2.0 Flash Experimental (Deepmind, 2025), Gemini 1.5 Pro (Team, 2024), and the open-source models Qwen2VL-72B-Instruct (Wang et al., 2024a), LLaVA-OneVision 72B (Li et al., 2024b) and InternVL 2.5 78B (Chen et al., 2024). For simplicity, in the following the models are referred to as o1, GPT-4o, Claude, Gemini 2.0 and 1.5, Qwen2VL, LLaVA-OneVision, and InternVL 2.5 respectively. The specific models and their configurations are given in Suppl.  A.2.
@@
When comparing BPs that have been solved at least once by any VLM to those solved at least once by a human, an interesting pattern emerges, as visualized in Figure 4 (right). Out of the $100100$ BPs, nearly all are solved by at least one human participant. In contrast, only $6666$ BPs are solved at least once by any model. A more detailed breakdown of the model’s results can be found in Figure 8 (right). The results show that there exist a group of BPs that remains particularly challenging for current state-of-the-art VLMs.
@@
Intuitively, one might expect that if a model can solve a BP in Task 1, it should also be able to solve Task 2 for this BP. To examine this, we compare the set of BPs solved in Task 1 to those successfully completed in Task 2. In Task 1, we define a BP as solved if it is correctly identified in at least two out of three attempts. Similarly, in Task 2, a BP is considered solved if all its images are correctly classified in at least two out of three attempts. We visualize these two sets and their intersection in Figure 10. Surprisingly, we find that the overlap between these two sets is quite small. This suggests that for a substantial subset of BPs, the models are able to solve them in Task 1 despite failing to classify every individual image correctly according to the ground truth rule. Conversely, there are also many BPs where the models perfectly classify images but still fail to solve the open-ended task. This large discrepancy highlights a surprising gap between recognizing correct classifications and effectively applying that knowledge in problem-solving. We report the number of correctly identified concepts for Task 2 in Table 8 (Suppl.).
For a better assessment of these results, we give a qualitative impression of the models’ mistakes when prompted to identify specific, rather simple, concepts of BP#16 and BP#55 in Figure 6 (cf. Table 12, 13 for all responses and Figure 9 for additional BP#29 and #37). We find that for BP#16 (cf. Figure 6 top left), even though some images are classified correctly, none of the models can classify all images correctly. Instead, we can see a tendency to classify one of the directions rather than the other, e.g., Claude and GPT-4o almost always predict clockwise as the turning direction for all of the diagrams.
@@
For BP#55 (cf. Figure 6 top right) we also see that none of the models is able to classify the concepts of all images correctly. Here, we can see very similar behaviors across the different models, e.g., almost all models categorize the first diagram of the BP falsely, but in most of the attempts, they are able to categorize the third diagram correctly. Interestingly, we found a strong correlation between the absolute position of the circle in the diagram (left vs. right) and the models responses. We take this as an indicator, that the models did not manage to reason spatially about the position of the circle in relation to the cavity.
@@
Further, we observe that for most models Task 1 is not a subset of Task 3. This is somewhat surprising, as one might expect that if a model can solve the BP, it would also include a correct rule in the set of hypotheses for Task 3. This finding may suggest a lack of robustness in handling the task.
@@ 5 Discussion: Remaining Challenges for AI
Most challenging BPs. Throughout the experiments, we saw that the investigated VLMs can solve some BPs better than others, where the set of solved puzzles appears to differ across task setting. Interestingly, we find that there are $1313$ BPs that are especially hard, i.e., they were not solved in any of the tasks (cf. Suppl.  B.1). We think that these BPs can serve as a good foundation for further research and bench-marking.
@@ 6 Conclusion and Future Work
Our findings raise several critical questions: Why do VLMs encounter difficulties with seemingly simple Bongard Problems, despite performing impressively across various established VLM benchmarks (Li et al., 2024a; Duan et al., 2024)? How meaningful are these benchmarks in assessing true reasoning capabilities? Future work could use our results as base to develop new benchmarks. Further, the discovered Type II subsets suggest opportunities for improving the performance, e.g., via a multi-stage reasoning approach.
@@ Appendix B Additional Results
In the following the detailed results of the evaluations are presented. In Table 4 the results of of each model on Task 1 are reported. Further, it includes the results of our human study. The detailed results for Task 1.1 are given in Table 5 (100 Options) and Table 6 (10 Options).
In Figure 8, we report detailed results of our evaluations regarding (Q2). We provide the results of each model type individually over each three runs and also the top 5 human study participants individually.
We provide quantitative results for Task 2 in Table 8 and discuss additional qualitative results for the evaluations of (Q3), i.e., also for BP#29 and #37. For BP#29, the models were asked to determine whether there are more shapes inside or outside the bigger shape (cf. Figure 6 bottom left). Even though the final decisions of the models were primarily correct, we saw in the answers that except for Claude, none of the models was able to count the shapes correctly.
We provide a full visualisation of all results over all tasks and including the different behavior types in Figure 10.
@@
Table 7: Percentages of solved BPs for humans and VLMs across categories.
@@
Task 2 Task 2 $||$ T1 $\cap$ T2 $||$ $||$ T1 $\cap$ T2 $||$ / T1
Model All 3/3 All 2/3 (Both 2/3) (Both 2/3)
o1 24 24 17 31.48%
GPT-4o 19 24 15 46.88%
Claude 16 22 15 34.88%
Gemini 2.0 Flash 19 30 18 45.00%
Gemini 1.5 Pro 20 22 4 36.36%
LLaVA-OneVision 11 15 4 33.33%
Qwen2VL 20 20 6 37.50%
InternVL2.5 10 16 10 34.48%
@@
For Task 2 we report the classification accuracy of each model for the ground truth concepts of each BP in Table 9. Further, we report the classification results of the qualitative examples of BP#16 (Table 12), BP#29 (Table 13, BP#36 (Table 14) and BP#55 (Table 15).

### v4

@@ 1 Introduction
Visual reasoning, the ability to understand, interpret, and reason about the visual world, is a fundamental aspect of human intelligence (Marr, 2010). It allows us to navigate our environment, interact with objects, and make sense of complex visual scenes. In recent years, the field of artificial intelligence (AI) has advanced rapidly toward replicating aspects of this visual reasoning, with significant focus placed on Vision-Language Models (VLMs) (Bordes et al., 2024; Liu et al., 2023, 2024a). These models integrate visual and textual information to generate descriptive content, aiming to mimic how humans comprehend and reason about the physical world. Because of their human-like responses, VLMs often create the illusion of possessing human-like perception and intelligence. However, as recent work shows, VLMs and the Large Language Models (LLM) on which they are based have dramatic shortcomings in the case of reasoning (Nezhurina et al., 2024) and visual perception (Kamath et al., 2023; Rahmanzadehgervi et al., 2024; Geigle et al., 2024; Gou et al., 2024) or their combination (Zhou et al., 2023; Zhang et al., 2024; Wang et al., 2024b).
@@
While traditional machine learning approaches have made some early progress on BPs (Raghuraman et al., 2023; Depeweg et al., 2024), the potential of VLMs remains largely unexplored. Given that VLMs already struggle with recognizing relatively simple visual patterns (Rahmanzadehgervi et al., 2024; Zhang et al., 2024), BPs present a particularly challenging task, offering a valuable framework for investigating which patterns are easier or harder for state-of-the-art models to identify. Recent work has explored BPs in the context of VLM evaluation (Małkiński et al., 2025). While their study provides meaningful insights, it does not analyze model behavior and failures in greater depth. This highlights the need for a more comprehensive investigation into how VLMs process and reason about BPs.
@@ 2 Related Work
Benchmarks for VLMs. Benchmarks specifically designed for VLMs usually involve complex tasks such as image captioning, scene or diagram understanding, visual question answering (VQA), or visual-commonsense reasoning (Masry et al., 2022; Yue et al., 2024; Rahmanzadehgervi et al., 2024; Kamath et al., 2023; Antol et al., 2015; Johnson et al., 2017; Hong et al., 2021; Zellers et al., 2019; Hudson & Manning, 2019; Qiao et al., 2024; Liu et al., 2024b). Yet, most of these only require simple reasoning abilities. More recent benchmarks have been introduced to probe advanced reasoning skills, e.g., logical learning (Helff et al., 2025; Vedantam et al., 2021; Xiao et al., 2024), mathematical reasoning (Lu et al., 2024) or analogy-based visual reasoning (Chollet, 2019; Moskvichev et al., 2023; Zhang et al., 2019). Małkiński et al., 2025 considered BPs for evaluating VLMs, concentrating on open-ended and classification-based settings. Our work shares their open-ended focus but goes further by adding two additional tasks to examine specific abilities we hypothesize are vital for solving Bongard problems. While Małkiński et al., 2025 compare synthetic Bongard problems to a newly proposed real-world variant, we target the original Bongard set, pinpointing especially challenging cases and identifying concept-detection inconsistencies. Generally, the shift towards more cognitively demanding tasks is promising, comprehensive diagnostic evaluations of VLMs’ reasoning capabilities that pinpoint sources of error and model limitations remain scarce. Furthermore, the degree to which these models genuinely comprehend complex, abstract visual concepts is yet to be fully investigated.
Open-Ended Visual Reasoning. Most existing benchmarks rely on classification tasks (Chollet, 2019; Helff et al., 2025; Vedantam et al., 2021), multiple-choice formats (Zhang et al., 2019; Lu et al., 2024), or synthetic environments with a limited set of operations (Moskvichev et al., 2023). While these setups tackle pattern recognition and visual reasoning, they do not fully capture real-world complexity. Open-ended visual puzzles, such as BPs, present a greater challenge by requiring models to map visual input to text without predefined rules or language biases. This lack of structured guidance makes the task more demanding, better reflecting real-world reasoning.
@@ 3 Bongard Problems and VLMs
To achieve this, we designed a tailored prompt for each BP that directly targets its ground truth concepts. For each of the 12 images in a BP, the prompt asks whether the concept from the left or right side of the problem applies to that specific image. To generate these prompts, we provide an LLM with example prompts for BP#16 and BP#55 (cf. Suppl.  A.1 and Suppl.  A.1) along with the corresponding ground truth rule [^2] https://www.foundalis.com/res/bps/bongard_problems_solutions.htm. Based on this input, the LLM automatically generates a custom prompt for each BP (cf. Suppl.  A.1).
@@ 4 Experimental Evaluation
Our experimental evaluations investigate to what extent state-of-the-art VLMs can solve Bongard problems. For this, we first evaluate the models quantitatively on all 100 puzzles based on the settings of Task 1 and Task 1.1. We then compare their performance against humans and investigate them qualitatively in more detail based on the settings of Task 2 and Task 3. We hereby address the following research questions:
  • (Q1)
    How well can state-of-the-art VLMs solve Bongard problems? (Task 1 + Task 1.1)
  • (Q2)
    How do VLMs perform in comparison to humans?
  • (Q3)
    How accurately can VLMs classify images of the BPs given the ground truth rule? (Task 2)
  • (Q4)
    How accurately can VLMs generate an underlying hypothesis? (Task 3)
@@
Models. We evaluated the proprietary models o1 (OpenAI, 2024b), GPT-4o (OpenAI, 2024a), Claude 3.5 Sonnet (Antropic, 2024), Gemini 2.0 Flash Experimental (Deepmind, 2025), Gemini 1.5 Pro (Team, 2024), and the open-source models Qwen2VL-72B-Instruct (Wang et al., 2024a), LLaVA-OneVision 72B (Li et al., 2025) and InternVL 2.5 78B (Chen et al., 2024). For simplicity, in the following the models are referred to as o1, GPT-4o, Claude, Gemini 2.0 and 1.5, Qwen2VL, LLaVA-OneVision, and InternVL 2.5 respectively. The specific models and their configurations are given in Suppl.  A.2.
@@
When comparing BPs that have been solved at least once by any VLM to those solved at least once by a human, an interesting pattern emerges, as visualized in Figure 4 (right). Out of the $100100$ BPs, nearly all are solved by at least one human participant. In contrast, only $6666$ BPs are solved at least once by any model. A more detailed breakdown of the model’s results can be found in Figure 8 (bottom). The results show that there exist a group of BPs that remains particularly challenging for current state-of-the-art VLMs.
@@
Intuitively, one might expect that if a model can solve a BP in Task 1, it should also be able to solve Task 2 for this BP. To examine this, we compare the set of BPs solved in Task 1 to those successfully completed in Task 2. In Task 1, we define a BP as solved if it is correctly identified in at least two out of three attempts. Similarly, in Task 2, a BP is considered solved if all its images are correctly classified in at least two out of three attempts. We visualize these two sets and their intersection in Figure 5. Surprisingly, we find that the overlap between these two sets is quite small. This suggests that for a substantial subset of BPs, the models are able to solve them in Task 1 despite failing to classify every individual image correctly according to the ground truth rule. Conversely, there are also many BPs where the models perfectly classify images but still fail to solve the open-ended task. This large discrepancy highlights a surprising gap between recognizing correct classifications and effectively applying that knowledge in problem-solving. We report the number of correctly identified concepts for Task 2 in Table 9 (Suppl.) and further investigate in Suppl.  B.4 whether the observed gap is affected by presenting full BP images versus individual images to the VLM.
@@
For a better assessment of these results, we give a qualitative impression of the models’ mistakes when prompted to identify specific, rather simple, concepts of BP#16 and BP#55 in Figure 6 (cf. Table 10, 11 for all responses and Figure 9 for additional BP#29 and #37). We find that for BP#16 (cf. Figure 6 top), even though some images are classified correctly, none of the models can classify all images correctly. Instead, we can see a tendency to classify one of the directions rather than the other, e.g., Claude and GPT-4o almost always predict clockwise as the turning direction for all of the diagrams.
For BP#55 (cf. Figure 6 bottom) we also see that none of the models is able to classify the concepts of all images correctly. Here, we can see very similar behaviors across the different models, e.g., almost all models categorize the first diagram of the BP falsely, but in most of the attempts, they are able to categorize the third diagram correctly. Interestingly, we found a strong correlation between the absolute position of the circle in the diagram (left vs. right) and the models responses. We take this as an indicator, that the models did not manage to reason spatially about the position of the circle in relation to the cavity.
@@
Furthermore, we observe that for most models, the set of BPs solved in Task 1 is not a subset of those with correctly identified rules in Task 3. This is somewhat unexpected, as one might assume that successfully solving a BP would imply the correct rule is among the hypotheses generated in Task 3. This finding may point to a lack of robustness in generalizing across the different task setups.
@@ 5 Discussion: Remaining Challenges for AI
Most challenging BPs. Throughout the experiments, we saw that the investigated VLMs can solve some BPs better than others, where the set of solved puzzles appears to differ across task setting. Interestingly, we find that there are $1010$ BPs that are especially hard, i.e., they were not solved in any of the tasks (cf. Suppl.  B.7). We think that these BPs can serve as a good foundation for further research and bench-marking.
@@ 6 Conclusion and Future Work
Our findings raise several critical questions: Why do VLMs encounter difficulties with seemingly simple Bongard Problems, despite performing impressively across various established VLM benchmarks (Li et al., 2024; Duan et al., 2024)? How meaningful are these benchmarks in assessing true reasoning capabilities? Future work could use our results as base to develop new benchmarks. Further, the discovered Type II subsets suggest opportunities for improving the performance, e.g., via a multi-stage reasoning approach.
Acknowledgements
@@ Appendix B Additional Results
In the following the detailed results of the evaluations are presented.
B.1 Task 1 and 1.1
In Table 4 the results of of each model on Task 1 are reported. Further, it includes the results of our human study. The detailed results for Task 1.1 are given in Table 5 (100 Options) and Table 6 (10 Options).
@@
In Figure 8, we report detailed results of our evaluations regarding (Q2). We provide the results of each model type individually over each three runs and also the top 5 human study participants individually. Corresponding numerical results are provided in Table 7.
Table 7: Mean percentages of solved BPs for humans and VLMs across categories.
@@
B.3 Task 2
For Task 2 we report the classification accuracy of each model for the ground truth concepts of each BP in Table 8. We provide quantitative results for Task 2 visually in Table 9 and discuss additional qualitative results for the evaluations of (Q3), i.e., also for BP#29 and #37. For BP#29, the models were asked to determine whether there are more shapes inside or outside the bigger shape (cf. Figure 9 bottom left). Even though the final decisions of the models were primarily correct, we saw in the answers that except for Claude, none of the models was able to count the shapes correctly.
@@ B.3 Task 2
Task 2 Task 2 $||$ T1 $\cap$ T2 $||$ $||$ T1 $\cap$ T2 $||$ / T1
Model All 3/3 All 2/3 (Both 2/3) (Both 2/3)
o1 24 24 17 31.48%
GPT-4o 19 24 15 46.88%
Claude 16 22 15 34.88%
Gemini 2.0 Flash 19 30 18 45.00%
Gemini 1.5 Pro 20 22 4 36.36%
LLaVA-OneVision 11 15 4 33.33%
Qwen2VL 20 20 6 37.50%
InternVL2.5 10 16 10 34.48%
For completeness, we report image-wise classification results of the qualitative examples of BP#16 (Table 10), BP#29 (Table 11, BP#37 (Table 12) and BP#55 (Table 13).
Clockwise Counter-clockwise
1 2 3 4 5 6 7 8 9 10 11 12
o1 1/1 0/1 0/1 0/1 1/1 1/1 1/1 0/1 1/1 0/1 1/1 0/1
GPT-4o 3/3 3/3 3/3 3/3 3/3 3/3 0/3 0/3 0/3 0/3 0/3 0/3
Claude 3.5 2/3 3/3 1/3 2/3 3/3 2/3 0/3 1/3 0/3 1/3 0/3 1/3
Gemini 2.0 2/3 3/3 2/3 3/3 3/3 3/3 0/3 2/3 0/3 0/3 1/3 0/3
Gemini 1.5 3/3 3/3 2/3 2/3 3/3 3/3 0/3 0/3 0/3 0/3 2/3 0/3
Llava-Onevision 2/3 2/3 2/3 2/3 1/3 2/3 3/3 2/3 1/3 3/3 1/3 2/3
Qwen2VL 3/3 3/3 3/3 3/3 3/3 3/3 0/3 0/3 0/3 0/3 0/3 0/3
InternVL2.5 2/3 2/3 1/3 1/3 2/3 1/3 3/3 1/3 2/3 1/3 1/3 1/3
Inside Outside
1 2 3 4 5 6 7 8 9 10 11 12
o1 1/1 1/1 1/1 1/1 1/1 1/1 1/1 1/1 0/1 1/1 1/1 1/1
GPT-4o 3/3 3/3 3/3 3/3 3/3 3/3 3/3 0/3 0/3 1/3 2/3 3/3
Claude 3.5 3/3 3/3 3/3 3/3 3/3 3/3 3/3 3/3 3/3 3/3 3/3 3/3
Gemini 2.0 3/3 3/3 3/3 3/3 3/3 3/3 3/3 3/3 1/3 3/3 3/3 3/3
Gemini 1.5 3/3 3/3 3/3 3/3 3/3 0/3 3/3 3/3 3/3 3/3 3/3 3/3
Llava-Onevision 3/3 3/3 2/3 3/3 3/3 3/3 3/3 2/3 1/3 3/3 3/3 3/3
Qwen2VL 3/3 3/3 3/3 3/3 3/3 3/3 3/3 3/3 0/3 3/3 3/3 3/3
InternVL2.5 3/3 3/3 1/3 0/3 2/3 2/3 2/3 2/3 3/3 3/3 3/3 3/3
Triangle Circle
1 2 3 4 5 6 7 8 9 10 11 12
o1 1/1 1/1 1/1 1/1 1/1 1/1 1/1 1/1 1/1 1/1 1/1 1/1
GPT-4o 3/3 3/3 3/3 3/3 3/3 3/3 3/3 3/3 3/3 3/3 3/3 3/3
Claude 3.5 3/3 3/3 3/3 3/3 3/3 3/3 3/3 3/3 3/3 3/3 3/3 0/3
Gemini 2.0 3/3 3/3 3/3 3/3 3/3 3/3 3/3 3/3 3/3 3/3 3/3 2/3
Gemini 1.5 3/3 3/3 3/3 3/3 3/3 3/3 3/3 3/3 3/3 3/3 3/3 3/3
Llava-Onevision 3/3 3/3 3/3 3/3 3/3 3/3 3/3 3/3 0/3 3/3 3/3 3/3
Qwen2VL 3/3 3/3 3/3 3/3 3/3 3/3 3/3 3/3 0/3 3/3 3/3 3/3
InternVL2.5 3/3 2/3 3/3 2/3 3/3 3/3 3/3 3/3 1/3 3/3 3/3 3/3
Left Right
1 2 3 4 5 6 7 8 9 10 11 12
o1 1/1 1/1 1/1 1/1 0/1 0/1 1/1 0/1 0/1 0/1 0/1 1/1
GPT-4o 0/3 3/3 3/3 3/3 0/3 0/3 3/3 0/3 0/3 3/3 0/3 3/3
Claude 3.5 1/3 2/3 2/3 2/3 0/3 0/3 2/3 0/3 2/3 3/3 1/3 3/3
Gemini 2.0 0/3 3/3 3/3 3/3 0/3 0/3 3/3 0/3 0/3 0/3 0/3 3/3
Gemini 1.5 0/3 3/3 3/3 3/3 0/3 0/3 3/3 0/3 0/3 0/3 0/3 3/3
Llava-Onevision 0/3 3/3 3/3 3/3 0/3 0/3 3/3 0/3 0/3 3/3 0/3 3/3
Qwen2VL 0/3 3/3 3/3 3/3 0/3 0/3 3/3 0/3 0/3 3/3 0/3 3/3
InternVL2.5 0/3 3/3 1/3 3/3 0/3 0/3 3/3 0/3 2/3 1/3 0/3 2/3
B.4 Full vs. Single Images
To examine whether the gap between solving the BP in Task 1 and correctly classifying the underlying concepts in Task 2 arises from how the problem is presented (i.e., as a full image), we conduct an additional analysis using two of the top-performing models from Task 1. This time, we provide the models with the twelve individual images from the BP instead of a single composite image. The prompt remains largely unchanged, with only minor adjustments to the image introduction. We report the results in Table 14. The findings show that this change in setup does not significantly affect the number of correctly solved problems per task, nor does it reduce the observed gap between Task 1 and Task 2.
Table 14: Comparison of model performance on Task 1 and 2 with original setup (full image) and single images as input.
T1 w/o T2 T2 T2 w/o T1
GPT-4o (orig) 13 11 13
GPT-4o (single images) 12 12 12
Claude (orig) 20 11 11
Claude (single images) 20 12 10
B.5 Error Types
We provide a full visualization of all results over all tasks and including the different behavior types in Figure 10.
B.6 Task 3